OpenStudy (usukidoll):

constant coefficient problem y'''-2y''-y'+2y = 0

OpenStudy (usukidoll):

\[r^3-2r^2-r+2 = 0\] \[-r^2(-r+2)-r+2=0\] \[(-r+2)(-r^2+1)=0\] \[r = 1,-1,2\] \[c_1e^x+c_2e^{2x}+c_3e^{-x}\]

OpenStudy (usukidoll):



OpenStudy (usukidoll):